Find five consecutive even integers is the sum of the first and fifth is to less than three times the fourth
Answer by Boreal(15235)

I assume 2 less (fewer)
x, x+2, x+4,x+6,x+8
x+x+8+2=3(x+6)
2x+10=3x+18
x=-8
-8,-6,-4,-2,0
sum of -8 and 0=-8
that is 2 fewer than 3*(-2)